OpenStudy (anonymous):

what is the domain and range of y=1/(x-2)^2? (in terms of infinity).

OpenStudy (anonymous):

\[y=1\div(x-2)^{2}\]

OpenStudy (anonymous):

what is the domain and range of the equation? @lucaz

OpenStudy (lucaz):

you need to find the values of x that dosent make (x-2)^2 = 0

OpenStudy (anonymous):

what do you mean? @lucaz

OpenStudy (lucaz):

you have a fraction 1/(x-2)^2 the denominator cannot be 0, only positive or negative

OpenStudy (lucaz):

if you set x=1 it would be (1-2)^2=1 if you set x=4 it would be (4-2)^2=4 but if x=2 (2-2)^2=0 x cannot be equal to 2 the domain is all real numbers except 2

OpenStudy (lucaz):

got it?

OpenStudy (anonymous):

so would it be \[[2,\infty)\]?

OpenStudy (anonymous):

but then what would the range be?

OpenStudy (lucaz):

no, it's R\[(x \in \mathbb{R}\2)\]

OpenStudy (lucaz):

I think "except" is like that, I don't remember

OpenStudy (anonymous):

it means the same thing because it's saying that you can all real numbers to infinity but the "[" means it cannot be 2.

OpenStudy (anonymous):

at least thats what my teacher is teaching us in PreCalc.

OpenStudy (lucaz):

[2,∞) this means x>=2

OpenStudy (lucaz):

but you can pick a negative value too

OpenStudy (lucaz):

and a number less than 2

OpenStudy (lucaz):

whatever you want, except 2, because the denominator would be 0

OpenStudy (anonymous):

so would it be (2,∞) meaning the restriction of 2.

OpenStudy (lucaz):



OpenStudy (lucaz):

this is the domain

OpenStudy (anonymous):

and the range?

OpenStudy (lucaz):

I forgot how to write it, I'm trying to remember x__X

OpenStudy (lucaz):

ops, wrong

OpenStudy (lucaz):

I made a mistake

OpenStudy (lucaz):

the range is only greater than 0 [0,infinity)

OpenStudy (lucaz):

(0,infinity), sorry